A 6 pole generator has a lap wound armature with 40 slots with 20 conducters per slot. The Flux per pole is 25mWb. The terminal voltage is 288v and the voltage regulation of the generator is 4%. Determine the speed of the generator


1
Expert's answer
2021-05-07T08:12:30-0400

"E_g=\frac{P\phi NZ}{60A}"


"E_g=V+0.04V=288+0.04\cdot288=299.52\ (V)"


"N=\frac{60AE_g}{P\phi Z}=\frac{60\cdot6\cdot 299.52}{6\cdot0.025\cdot (40\cdot20)}=682\ (rpm)" . Answer